# ArMeme: Propagandistic Content in Arabic Memes

### v1

@@ 4.2.4 Multimodal LLMs
For the multimodal models Xu et al. (2023), we experimented with several well-known and top-performing commercial models. These included OpenAI’s GPT models (GPT-4 Turbo and GPT-4o) OpenAI (2023), as well as Google’s Gemini Pro models (versions 1.0 and 1.5) Team et al. (2023).
Using these models, we tested (i) the meme/image only, (ii) text only (text extracted using OCR from the image), and (iii) multimodal (meme and OCR text) in a zero-shot learning setting. This means we did not provide any training examples within the prompts to the models.
We designed a prompt based on trial and error using the visual interfaces of OpenAI’s GPT-4 user interface. The prompt instructs the models to perform a deeper analysis of the image and any text that they can read within the image before answering whether the meme can be classified as spreading propaganda. Additionally, it requests the models to provide the output in a valid JSON format. For the experiments, we used the default parameters for each multimodal model.
@@ 4.3 Prompting Strategy
LLMs produce varied responses depending on the prompt design, which is a complex and iterative process that presents challenges due to the unknown representation of information within different LLMs. The instructions expressed in our prompts include English language with the input text content in Arabic.
As mentioned earlier we employed zero-shot prompting, providing natural language instructions that describe the task and specify the expected output. This approach enables the LLMs to construct a context that refines the inference space, yielding a more accurate output. In Listing 1, we provide an example of a zero-shot prompt, emphasizing the instructions and placeholders for both input and label. Along with the instruction we provide the labels to guide the LLMs and provide information on how the LLMs should present their output, aiming to eliminate the need for post-processing.
@@
payload = {
"model": "gpt-4o",
"messages": [
"role": "user",
"content": [
"type": "text",
"text": prompt
"type": "image_url",
"image_url": {
"url": f"data:image/jpeg;base64,{base64_image}"
"type": "text",
"text": Text extracted: + t_text
@@
5 Results and Discussion
In Table 3, we report the detailed classification results for different modalities and models. All models outperform the majority and random baselines. Among the text-based models, the fine-tuned Qarib model outperforms all other models, achieving the best results (0.690 weighted F1) across all modalities and models. AraBERT is the second-best fine-tuned model, with a weighted F1-score of 0.666 among the text-based models. The performance of multilingual transformer models is relatively worse than that of monolingual models.
For the image-based models, the fine-tuned ResNet50 shows the best result (0.673 weighted F1) among all other fine-tuned models and GPT-4o model. The performance of MobileNet (v2) and CNeXt + SVM rank as the second and third best among the fine-tuned models. The results of VGG16 and EfficientNet (b7) are almost similar.
For the multimodal models, the model trained with ConvNeXt + AraBERT + SVM shows the highest performance (0.659 weighted F1) among the multimodal LLMs. The performance of Gemini is significantly worse than that of the GPT-4 variants. GPT-4o demonstrates higher performance compared to GPT-4 Vision.
In our experiments all multimodal model are tested using zero-shot setting, therefore, such lower performance compared to the fine-tuned models are expected.
@@ 6 Additional Experiments
We further conducted experiments using the dataset released as part of the ArAIEval shared task 2 Hasanain et al. (2024), focusing on two labels: propaganda and not-propaganda. The dataset statistics are provided in Appendix A.2. The goal was to investigate model performance in a binary classification scenario.
Table 4 presents the competitive results of four multimodal models with image-only input: GPT-4o, GPT-4 Turbo, and Gemini Pro 1.0. Among these models, GPT-4o significantly outperforms the others and demonstrates the highest performance across all evaluated metrics, achieving an accuracy of 85.17%, a precision of 84.80, a recall of 85.17, and a weighted F1-score of 84.87. In comparison, GPT-4 Turbo lags behind GPT-4o in all metrics, with an accuracy of 76.44%, indicating a significant performance drop compared to GPT-4o. Gemini Pro 1.0 shows lower performance than the GPT-4 models, with an accuracy of 72.47%.
@@ 7 Conclusions and Future Work
In this study, we introduce a manually annotated dataset for detecting propaganda in Arabic memes. We have annotated $\sim$ 6K memes with four different categories, making it the first such resource for Arabic content. To facilitate future annotation efforts for this type of content, we developed annotation guidelines in both English and Arabic and are releasing them to the community. Our work provides an in-depth analysis of the dataset and includes extensive experiments focusing on different modalities and models, including pre-trained language models (PLMs), large language models (LLMs), and multimodal LLMs. Our results indicate that fine-tuned models significantly outperform LLMs.
In future work, we plan to extend the dataset with further annotations that include hateful, offensive, and propagandistic techniques.
@@ Acknowledgments
The contributions of F. Alam, M. Hasanain, and F. Ahmed were funded by the NPRP grant 14C-0916-210015, which is provided by the Qatar National Research Fund (a member of Qatar Foundation).
Appendix A Additional Dataset Details
A.1 Data Sources
In Table 5, we provide the sources of the dataset, number of groups and number of image we have collected.
Source # of Group # of Images
Facebook 19 5,453
Instagram 22 107,307
Pinterest - 11,369
Twitter - 5,369
Total 129,498
Table 5: Statistics of the initial data collection.
Figure 3: Keywords used to collect tweets.
A.2 ArAIEval Shared Task Dataset
Table 6 report the distribution of the dataset released as a part of the ArAIEval shared task. For our study we benchmarked this dataset using multimodal models.
Class labels Train Dev Test Total
Not propaganda 1,540 224 436 2,200
Propaganda 603 88 171 862
Total 2,143 312 607 3,062
Table 6: Distribution of dataset for ArAIEval shared task 2.
Appendix B Details of the experiments
@@
Appendix C Annotation Task
We designed the annotation instructions through careful analysis and discussion, followed by iterative refinements based on observations and input from the annotators based on the pilot annotation. Our annotation schema is structured into two phases.
C.1 Phases of Annotations
To ensure the quality of the annotation and facilitate the work of annotators, we conducted the annotation in two phases: (i) meme categorization and (ii) text editing. The first phase (see Section C.2) focuses primarily on categorization. In the second phase (see Section C.3), our goal is to edit the text only for memes labeled as propagandistic or not propagandistic. The motivation for editing the text for these categories is to further utilize them for other annotation tasks. For example, propagandistic memes can be further annotated with specific propagandistic techniques. In Figure , we illustrate the thought process of the meme annotation phases.
@@ C.1 Phases of Annotations
C.2 Meme Categorization
C.2.1 Definition of a Meme:
@@
    Text overlaid on image.
@@
    Mostly uses images of entities with facial expressions (human, animals, fictional characters), which are then used to construct meaning alongside the added text.
@@
    Uses an entity performing a certain action that might be used to construct meaning alongside the added text.
@@
    Uses an entity that represents an idea or culture, to construct meaning alongside the added text.
@@
    Mostly uses screenshots from movie scenes and dialogues with added comments, to create memes.
@@
C.2.2 Defining Propaganda:
Propaganda is any communication that deliberately misrepresents symbols and/or entities, appealing to emotions and prejudices while bypassing rational thought, to influence its audience toward a specific goal. Memes are created to be humorous; therefore, it is natural that they lack rational discussion. Instead, they use content to appeal to emotions and prejudices. For our task, we defined the following four categories and annotated the memes accordingly.
@@ C.2.2 Defining Propaganda:
Note: Memes might contain words that have an implicitly offensive meaning, or the use of offensive words may be aimed at social, religious, or political groups. In these cases, the meme does not fall under this criterion.
@@
C.3 Text Editing
The task is to edit the text to match the text shown in the image. The interface will show you the picture, alongside the text that is viewed in it. The text was extracted automatically, so it might contain errors. It might not reflect all you see in the picture. Some important guidelines to follow for editing the text are listed below:
@@ C.3 Text Editing
Appendix D Annotation Platform
D.1 Meme Categorization Task
In Figure 11, we provide a screenshot of the annotation platform for the meme categorization task. As shown in the figure, the platform displays the meme itself on the right, the extracted text on the left, a link to the annotation guidelines, and labels with buttons at the bottom for selecting a category for the meme. The task of the annotator was to label the meme as one of the below categories, according to the definitions detailed in the guideline (see Section C). To facilitate the work of annotators in the annotation process, we used the keywords ‘meme’ along with the labels ‘other’, ‘propaganda’, and ‘not-propaganda’.
@@ D.1 Meme Categorization Task
Given that the memes we collected were from different social media platforms, they may contain offensive content. Therefore, we added a note that some pictures may contain offensive content, and that we apologize for any inconvenience that such content may cause. We appreciate your contribution to this project which will minimize the spread of such harmful content on the internet.
To further guide the annotation process, we asked the annotators to follow the following steps.
@@
D.2 Text Editing Task
@@
In this phase, the task was to edit the text based on the guidelines discussed in Section C.3. In Figure 12, we provide a screenshot demonstrating the text extracted from OCR, an editable text box, and the original meme. The task was to edit the text to match it with the original meme.

### v2

@@ 4.2.4 Multimodal LLMs
For the multimodal models Xu et al. (2023), we experimented with several well-known and top-performing commercial models. These included OpenAI’s GPT models (GPT-4 Turbo and GPT-4o) OpenAI (2023), as well as Google’s Gemini Pro models (versions 1.0 and 1.5) Team et al. (2023). Using these models, we tested (i) the meme/image only, (ii) text only (text extracted using OCR from the image), and (iii) multimodal (meme and OCR text) in a zero-shot learning setting. This means we did not provide any training examples within the prompts to the models. We designed a prompt based on trial and error using the visual interfaces of OpenAI’s GPT-4 user interface. The prompt instructs the models to perform a deeper analysis of the image and any text that they can read within the image before answering whether the meme can be classified as spreading propaganda. Additionally, it requests the models to provide the output in a valid JSON format. For the experiments, we used the default parameters for each multimodal model.
@@ 4.3 Prompting Strategy
LLMs produce varied responses depending on the prompt design, which is a complex and iterative process that presents challenges due to the unknown representation of information within different LLMs. The instructions expressed in our prompts include English language with the input text content in Arabic. As mentioned earlier we employed zero-shot prompting, providing natural language instructions that describe the task and specify the expected output. This approach enables the LLMs to construct a context that refines the inference space, yielding a more accurate output. In Listing 1, we provide an example of a zero-shot prompt, emphasizing the instructions and placeholders for both input and label. Along with the instruction we provide the labels to guide the LLMs and provide information on how the LLMs should present their output, aiming to eliminate the need for post-processing.
@@
5 Results and Discussion
In Table 4, we report the detailed classification results for different modalities and models. All models outperform the majority and random baselines. Among the text-based models, the fine-tuned Qarib model outperforms all other models, achieving the best results (0.690 weighted F1) across all modalities and models. AraBERT is the second-best fine-tuned model, with a weighted F1-score of 0.666 among the text-based models. The performance of multilingual transformer models is relatively worse than that of monolingual models. For the image-based models, the fine-tuned ResNet50 shows the best result (0.673 weighted F1) among all other fine-tuned models and GPT-4o model. The performance of MobileNet (v2) and CNeXt + SVM rank as the second and third best among the fine-tuned models. The results of VGG16 and EfficientNet (b7) are almost similar. For the multimodal models, the model trained with ConvNeXt + AraBERT + SVM shows the highest performance (0.659 weighted F1) among the multimodal LLMs. The performance of Gemini is significantly worse than that of the GPT-4 variants. GPT-4o demonstrates higher performance compared to GPT-4 Vision. In our experiments all multimodal model are tested using zero-shot setting, therefore, such lower performance compared to the fine-tuned models are expected.
@@ 6 Additional Experiments
We further conducted experiments using the dataset released as part of the ArAIEval shared task 2 Hasanain et al. (2024b), focusing on two labels: propaganda and not-propaganda. The dataset statistics are provided in Table 5. The goal was to investigate model performance in a binary classification scenario and we benchmarked this dataset using multimodal models.
Class labels Train Dev Test Total
Not propaganda 1,540 224 436 2,200
Propaganda 603 88 171 862
Total 2,143 312 607 3,062
Table 5: Distribution of dataset for ArAIEval shared task 2.
Table 6 presents the competitive results of three multimodal models with image-only input: GPT-4o, GPT-4 Turbo, and Gemini Pro 1.0. Among these models, GPT-4o significantly outperforms the others and demonstrates the highest performance across all evaluated metrics, achieving an accuracy of 85.17%, a precision of 84.80, a recall of 85.17, and a weighted F1-score of 84.87. In comparison, GPT-4 Turbo lags behind GPT-4o in all metrics, with an accuracy of 76.44%, indicating a significant performance drop compared to GPT-4o. Gemini Pro 1.0 shows lower performance than the GPT-4 models, with an accuracy of 72.47%.
@@ 7 Conclusions and Future Work
In this study, we introduce a manually annotated dataset for detecting propaganda in Arabic memes. We have annotated $\sim$ 6K memes with four different categories, making it the first such resource for Arabic content. To facilitate future annotation efforts for this type of content, we developed annotation guidelines in both English and Arabic and are releasing them to the community. Our work provides an in-depth analysis of the dataset and includes extensive experiments focusing on different modalities and models, including pre-trained language models (PLMs), large language models (LLMs), and multimodal LLMs. Our results indicate that fine-tuned models significantly outperform LLMs. In future work, we plan to extend the dataset with further annotations that include hateful, offensive, and propagandistic techniques.
@@ Acknowledgments
The work of F. Alam, M. Hasanain, and F. Ahmed is supported by the NPRP grant 14C-0916-210015 from the Qatar National Research Fund part of Qatar Research Development and Innovation Council (QRDI). The findings achieved herein are solely the responsibility of the authors.
Appendix A Details of The Experiments
@@
Appendix B Annotation Task
We designed the annotation instructions through careful analysis and discussion, followed by iterative refinements based on observations and input from the annotators based on the pilot annotation. Our annotation schema is structured into two phases as discussed below.
B.1 Phases of Annotations
To ensure the quality of the annotation and facilitate the work of annotators, we conducted the annotation in two phases: (i) image categorization and (ii) text editing. The first phase (see Section B.2) focuses primarily on categorizing the images shown on the interface. In the second phase (see Section B.3), our goal is to edit the text that can be seen on the images only for images that were labelled as memes and as propagandistic or not propagandistic. The motivation for editing the text for these categories is to further utilize them for other annotation tasks. For example, propagandistic memes can be further annotated with specific propagandistic techniques. In Figure 4, we illustrate the thought process of the meme annotation phases.
@@ B.1 Phases of Annotations
B.2 Meme Categorization
B.2.1 Definition of a Meme:
@@
    contains text overlaid on image.
@@
    May use images of entities with facial expressions (human, animals, fictional characters,etc.), which are then used to construct meaning alongside the added text.
@@
    May use an entity performing a certain action that might be used to construct meaning alongside the added text.
@@
    May use an entity that represents an idea or culture, to construct meaning alongside the added text.
@@
    May use screenshots from movie scenes and dialogues with added comments to create memes.
@@
B.2.2 Defining Propaganda:
Propaganda is any communication that deliberately misrepresents symbols and/or entities, appealing to emotions and prejudices while bypassing rational thought, to influence its audience toward a specific goal. Memes are created to be humorous; therefore, it is natural that they lack rational discussion. Instead, they use content to appeal to emotions and prejudices. For our task, we defined the following four categories and annotated the images accordingly.
@@ B.2.2 Defining Propaganda:
Note: Memes might contain words that have an implicitly offensive meaning, or uses offensive words may be aimed at social, religious, or political groups. In these cases, the meme does not fall under this criterion.
@@
B.3 Text Editing
The task is to edit the text to match the text shown in the image. The interface will show the picture, alongside the text that is in it. The text was extracted automatically, so it might contain errors. It might not reflect all that is sees in the picture. Some important guidelines to follow for editing the text are listed below:
@@ B.3 Text Editing
Appendix C Annotation Platform
C.1 Meme Categorization Task
In Figure 11, we provide a screenshot of the annotation platform for the meme categorization task. As shown in the figure, the platform displays the meme itself on the right, the extracted text on the left, a link to the annotation guidelines, and labels with buttons at the bottom for selecting a category for the meme. The task of the annotator was to label the meme as one of the below categories, according to the definitions detailed in the guideline (see Section B). To facilitate the work of annotators in the annotation process, we used the keywords ‘meme’ along with the labels ‘other’, ‘propaganda’, and ‘not-propaganda’.
@@ C.1 Meme Categorization Task
Given that the memes we collected were from different social media platforms, they may contain offensive content. Therefore, we added a note that some pictures may contain offensive content, and that we apologize for any inconvenience that such content may cause. We appreciate your contribution to this project which will minimize the spread of such harmful content on the internet. To further guide the annotation process, we asked the annotators to follow the following steps.
@@
C.2 Text Editing Task
@@
In this phase, the task was to edit the text based on the guidelines discussed in Section B.3. In Figure 12, we provide a screenshot demonstrating the text extracted from OCR, an editable text box, and the original meme. The task was to edit the text to match it with the original meme.
Appendix D Arabic Annotation Guideline
D.1 Meme Categorization
D.1.1 Definition of Meme
{RLtext}
"الميم" أو "الميمز" نص يتكون من صورة تمثل خلفية يضاف إليها نصٌ مكتوبٌ كتعليق أو مزحة، ويمكن لصورة الخلفية أن تكون صورة فوتوغرافية أو رسم توضيحي أو لقطة شاشة، ومن الشائع أن يتم وضع النص في أعلى أو أسفل الصورة ولكن ليس دائمًا، ويؤدي الجمع بين الصورة والنص إلى إنشاء رسالة أو نكتة أو تعليق معين من المفترض أن يسهل فهمه، والتأثر به، ويسهل نشره. متى ستكون الصورة المعروضة "ميمًا"؟
  1. 1.
    {RLtext}
    تجمع بين نص مكتوب وصورة
  2. 2.
    {RLtext}
    النص المكتوب هو تعليق فكاهي
  3. 3.
    {RLtext}
    لابد للصورة أن تستوفي الشرطين بالإضافة إلى بعض الخصائص المذكورة بالأسفل.
  4. 4.
    {RLtext}
    التعديل على الصورة المستخدمة في الميم.
  5. 5.
    {RLtext}
    إضافة نص مكتوب على أماكن متفرقة على الصورة.
  6. 6.
    {RLtext}
    يكثر استخدام صور تعبيرية، أي أنه في وجود كيان (إنسان، حيوان، شخصيات خيالية) بوجه في الصورة، يستخدم التعبيرات المرسومة على الوجه لإنشاء معنى مع التعليق المكتوب.
  7. 7.
    {RLtext}
    قد يكون في الصورة كيان يقوم بفعل ما، فيستخدم الفعل لإنشاء معنى مع التعليق المكتوب.
  8. 8.
    {RLtext}
    قد يكون في الصورة كيان (منظمة أو شخص)، فيستخدم ما يمثله ذلك الكيان لإنشاء معنى مع التعليق المكتوب.
  9. 9.
    {RLtext}
    يكثر استخدام صور وحوارات مأخوذة من أفلام ومسلسلات وإضافة تعليق عليها.
  10. 10.
    {RLtext}
    تتميز معظم الصور بأنه يمكن إعادة استخدامها بإضافة تعليق فكاهي آخر.
{RLtext}
ملاحظة: إنشاء المعنى مختلف عن إضافة المعنى، فنقصد في النقاط 6-8 أن عدم وجود تعبير الوجه أو الفعل أو الكيان سيغير المعنى أو سيكون المعنى غير مكتمل.
D.1.2 Definition of Propaganda
{RLtext}
في حكمك على "الميم" من حيث استخدامها لأسلوب من أساليب البروباغاندا، ستستند على التعريف التالي للبروباغاندا: شكل من أشكال التواصل يقوم فيه صاحب الرسالة – متعمدًا – بتشويه الرموز أو إثارة المشاعر أو التحيزات دون اللجوء إلى حجج منطقية وذلك للتأثير على الجمهور ودفعهم نحو هدف معين. بما أن "الميم" هو نص في أصله فكاهي، فعلى الأرجح أنك لن تجد فيها حجج منطقية، بل سيستخدم النص (الصورة أو النص المكتوب أو الإثنين) السخرية لتشويه الرموز وإثارة المشاعر.
(1) Not-Meme:
Figure 13 {RLtext} يندرج تحت هذا التصنيف الصور التي لا تتبع تعريف "الميمز" المذكور في هذا الدليل.
Figure 13: Examples of images labeled as not-meme.
(2) Other: Figure 14 {RLtext} يندرج تحت هذا التصنيف الصور التي تتبع تعريف "الميمز" ولكنها تقع تحت واحدة من المعايير المُحددة في القسم أدناه. المعايير والأمثلة:
  1. 1.
    {RLtext}
    "الميمز" التي تعتمد على مناظر وشخصيات كرتونية، ما عدا التي تحتوي على معنى مسيء لشخصيات سياسية أو مشهورة أو جماعات وأحزاب معينة.
  2. 2.
    {RLtext}
    "الميمز" التي تعتمد بشكل كلي على أرقام أو مخططات أو رسوم بيانية.
  3. 3.
    {RLtext}
    "الميمز" التي تظهر عري صريح.
  4. 4.
    {RLtext}
    "الميمز" التي تستخدم كلمات نابية ومنحطة وخادشة للحياء بشكل صريح (مثال: ابن الـ******).
  5. 5.
    {RLtext}
    "الميمز" التي تستخدم لغة ثانية غير العربية. (لا يوجد مثال)
  6. 6.
    {RLtext}
    "الميمز" التي لم تتمكن من فهمها بسبب اللهجة أو حجم الخط أو لأي سبب آخر. (لا يوجد مثال)
{RLtext}
قد تحتوي "الميمز" على كلمات منحطة لكن بشكل ضمني، أو على شتيمة صريحة غرضها تشويه سمعة أو التقليل (مثال: المنافق، السياسي الفاسد). لا تندرج هذه الكلمات أو العبارات تحت هذا التصنيف، فاختر لها ما يناسبها من التصنيفات الأخرى.
Figure 14: Examples of images labeled as other.
(3) Not Propaganda:
Figure 15 {RLtext} يندرج تحت هذا التصنيف الصور التي تتبع تعريف "الميمز" ولكنها لا تحتوي على بروباغاندا.
Figure 15: Examples of images labeled as not propaganda.
Figure 16: Examples of images labeled as propaganda.
(4) Propaganda:
Figure 16 {RLtext} يندرج تحت هذا التصنيف الصور التي تتبع تعريف "الميمز" وتحتوي على بروباغاندا.
D.2 Text Editing
{RLtext}
المطلوب منك في هذه المهمة هو تحرير نصوص عربية، وتعديلها لتطابق المذكور في الصورة المعروضة. ستعرض لك الواجهة صورة مع النص المستخرج منها آليًا، وقد يتضمن النص المستخرج أخطاء أو قد يكون ناقصًا، لكن لا نقصد هنا أخطاء إملائية أو أخطاء نحوية، بل نقصد أن النص لا يطابق المذكور في الصورة. إليك بعض الإرشادات المهمة:
  1. 1.
    {RLtext}
    يجب كتابة كل جملة مستقلة تشكّل معنى في سطر واحد.
  2. 2.
    {RLtext}
    علامات الترقيم الواضحة في الصورة تعتبر جزءًا من النص. يجب تحريرها أو إضافتها حسب الحاجة.
  3. 3.
    {RLtext}
    إذا كان النص في الصورة معروض في أعمدة، ضع أولاً كل النصوص في العمود الأول، ثم كل النصوص في العمود التالي. ستعمل على نصوص عربية لذا يجب اعتبار العمود الأول من اليمين، لكن الأمر متروك لفهمك لمعنى النص، المهم أن ترتب الجمل لتعطي المعنى المطلوب.
  4. 4.
    {RLtext}
    قم بإعادة ترتيب النص بحيث يكون هناك جملة واحدة في كل سطر، إذا كان ذلك ممكنًا.
  5. 5.
    {RLtext}
    إذا كانت هناك كتل أو أجزاء نصية منفصلة في الصورة، ابدأ سطرًا جديدًا لكل كتلة.
  6. 6.
    {RLtext}
    اترك فراغًا بين كتلتين من النص إذا عرضا في موضعين مختلفين في الصورة.
  7. 7.
    {RLtext}
    أزل العناصر التالية من النص:
    • {RLtext}
      أسماء المستخدمين وأسماء حسابات وسائل التواصل الاجتماعي لو كانت ظاهرة في الصورة
    • {RLtext}
      أي روابط أو نصوص أو شعارات لا تعتبر جزءا من النص الذي يشكل المعنى.
    • {RLtext}
      أي نص مخفي ويصعب قراءته.
  8. 8.
    {RLtext}
    في حالات خاصة قد تجد أن شعارًا ما استخدم في "الميمز" لإنشاء معنى، عندها أضف نص الشعار إلى النص الذي تحرره.
Example 1:
Figure 17 {RLtext} ملاحظات على تعديل النص في الصورة:
  1. 1.
    {RLtext}
    كل مربع حوار يعتبر جملة واحدة
  2. 2.
    {RLtext}
    إبدأ سطرًا جديدًا لكل مربع (كل مربع هو كتلة نصية مختلفة)
  3. 3.
    {RLtext}
    أزل أي عناصر لا تشكل جزءًا من المعنى: اسم الحساب والموقع
  4. 4.
    {RLtext}
    ضف أو عدل علامات الترقيم لتناسب المعروض في النص
  5. 5.
    {RLtext}
    النص بعد التعديل:
{RLtext}
شد….شد احصرو…جيبو خلي نعطيه حقوقو ايجا…جيتشي؟؟ ايجا …خوذ حقوقك يا ولد لحرام والله كان نشدك لما…. نعطيك حقوقك كاملين يا فرخ …
Figure 17: An example of a meme for editing text.
Example 2:
Figure 18 {RLtext} ملاحظات على تعديل النص في الصورة:
  1. 1.
    {RLtext}
    يصعب فهم وقراءة النص المكتوب باللون الأحمر، لذا يجب عدم وضعه من ضمن النص
  2. 2.
    {RLtext}
    النص المكتوب على القبعة والنص الذي بالأسود كل منهم كتلة نصية مختلفة، إبدأ سطرًا لكل منهم واترك فراغًا بس كل سطر جديد.
  3. 3.
    {RLtext}
    هذا المثال لتوضيح حالة فقط، ولن يتم عرض "ميمز" بالإنجليزية في هذه المهمة.
  4. 4.
    {RLtext}
    النص بعد التعديل:
            Bernie
            Ridin with Biden **2020**
            Haha hey its the Obama guy
Figure 18: An example of a meme for editing text.
Example 3:
Figure 19 {RLtext} ملاحظات على تعديل النص في الصورة:
  1. 1.
    {RLtext}
    تشكل شعار الجامعة جزءا من المعنى لذا يجب إضافته إلى النص
  2. 2.
    {RLtext}
    نضيف فقط الجزء الذي يسهل قراءته
  3. 3.
    {RLtext}
[…]
Figure 19: An example of a meme for editing text.